$SPAD/src/input f04maf.input

The Axiom Team December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1994.

———-

— * —

\getchunk{license}
)cl all -- first run f01maf
showArrayValues true
showScalarValues true
n := 4;
m := 4;
nn := n*m;
licn := 90;
lirn := 50;
a := new(1,licn,0.0)$Matrix DoubleFloat;
irn := new(1,lirn,0)$Matrix Integer;
icn := new(1,licn,0)$Matrix Integer;
k := 0;
for i in 1..nn repeat
  k := k + 1
  a(1,k) := 1
  irn(1,k) := i
  icn(1,k) := i
for i in 1..n repeat
  for j in 1..(m-1) repeat
    k := k + 1
    jj := (i-1)*n + j
    a(1,k) := -0.25
    irn(1,k) := jj
    icn(1,k) := jj + 1
for i in (n+1)..nn repeat
  k := k + 1
  a(1,k) := -0.25
  irn(1,k) := i-n
  icn(1,k) := i
nz := k;
abort := [true,true,true];
droptl := 0.1;
densw := 0.8;
ifail := 111;
res1:=f01maf(nn,nz,licn,lirn,abort,a,irn,icn,droptl,densw,ifail)
a := res1."avals"
irn := res1."irn"
icn := res1."icn"

b := [[0.5,0.25,0.25,0.5,0.25,0,0,0.25,0.25,0,0,0.25,0.5,0.25,0.25,0.5]]
acc := [[0.0001,0]]
noits := [[50,0]]
wkeep := res1."wkeep"
ikeep := res1."ikeep"
inform := res1."inform"
result:=f04maf(nn,nz,a,licn,irn,lirn,icn,wkeep,ikeep,inform,b,acc,noits,ifail)

———-

References^{[1] nothing}

5